# Active Adaptive Experimental Design for Treatment Effect Estimation with Covariate Choices

### v1

@@ B.2 Parametric Submodels
$f(y,a,x;\beta)=\prod_{d\in\{1,0\}}\Big{\{}r^{d}(y\mid x;\beta)w(d\mid x;\beta)\Big{\}}^{a=d}p(x;\beta)$ (4)
@@
We also denote $\prod_{d\in\{1,0\}}\Big{\{}r^{d}(y\mid x;\beta)w(d\mid x;\beta)\Big{\}}^{a=d}$ by $f(y,a\mid x;\beta)$.
@@
$\sum^{T}_{t=1}\log\big{(}f(Y_{t},A_{t},X_{t};\beta)+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\big{)}.$
By taking the derivative of $\sum^{T}_{t=1}\log\big{(}f(Y_{t},A_{t},X_{t};\beta)+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\big{)}$ with respect to $\beta$, we can obtain the corresponding score as
$S(W;\beta)\coloneqq\frac{\mathrm{d}}{\mathrm{d}\beta}\left\{\sum^{T}_{t=1}\log\big{(}f(Y_{t},A_{t},X_{t};\beta)+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\big{)}\right\}$
$=\sum^{T}_{t=1}\left\{\mathbbm{1}[A_{t}=1]\left\{s^{1}(Y_{t}\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right\}+\mathbbm{1}[A_{t}=0]\left\{s^{0}(Y_{t}\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right\}+\zeta(X_{t};\beta)\right\}$
@@ B.3 ATE Function
B.4 Cramér-Rao Lower Bound
@@
$s^{d}(y\mid x;\beta_{0})$ $=\frac{1}{T}\frac{\big{(}y-\mathbb{E}[Y(d)\mid X=x]\big{)}}{w(d\mid x)}$ (5)
@@
$\zeta(x;\beta_{0})$ $=0,$
$\widetilde{\zeta}(\widetilde{x};\beta_{0})$ $=\frac{1}{S}\Big{\{}\theta_{0}(\widetilde{x})-\theta_{0}\Big{\}},$
where we define
$\theta_{0}(\widetilde{x})\coloneqq\mu_{0}(1)(\widetilde{x})-\mu_{0}(0)(\widetilde{x}).$
@@
$\mathbb{E}_{W}\left[\left(\nabla_{\beta}\left\{\sum^{T}_{t=1}\log\big{(}f(Y_{t},A_{t},X_{t};\beta)\big{)}+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\right\}\right)^{2}\right]^{-1}.$
@@
$\mathrm{CR}(\theta)\coloneqq\nabla_{\beta}\theta(\beta)\mathbb{E}_{W}\left[\left(\nabla_{\beta}\left\{\sum^{T}_{t=1}\log\big{(}f(Y_{t},A_{t},X_{t};\beta)\big{)}+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\right\}\right)^{2}\right]^{-1}\nabla_{\beta}\theta(\beta).$
Theorem B.2.
@@
$C$ $=\mathbb{E}_{W}\Big{[}Y(1)s^{1}(Y(1)\mid X;\beta_{0})-Y(0)s^{0}(Y(0)\mid X;\beta_{0})+\theta(\widetilde{X};\beta_{0})\widetilde{\zeta}(\widetilde{X};\beta_{0})\Big{]},$
$D$ $=\mathbb{E}_{W}\left[\left(\nabla_{\beta}\left\{\frac{1}{T}\sum^{T}_{t=1}\log\big{(}f(Y_{t},A_{t},X_{t};\beta_{0})\big{)}+\frac{1}{S}\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta_{0})\right\}\right)^{2}\right].$
@@
$T\mathrm{CR}\big{(}\theta\big{)}$ $=T\mathbb{E}_{W}\left[\frac{1}{T}\left\{Y(1)\frac{\big{(}Y(1)-\mathbb{E}[Y(1)\mid X=x]\big{)}}{w(1\mid X)}+Y(0)\frac{\big{(}Y(0)-\mathbb{E}[Y(0)\mid X=x]\big{)}}{w(0\mid X)}\right\}+\frac{1}{S}\theta^{2}(\widetilde{X};\beta_{0})\right]^{2}$
×$\ \ \ \times\mathbb{E}_{W}\left[T\left\{\nabla_{\beta}\log\big{(}f(Y,A,X;\beta_{0})\big{)}\right\}^{2}+S\left\{\nabla_{\beta}\log\big{(}q(\widetilde{X};\beta_{0})\right\}^{2}\right]^{-1}$
$=\mathbb{E}_{W}\left[\left\{Y(1)\frac{\big{(}Y(1)-\mathbb{E}[Y(1)\mid X=x]\big{)}}{w(1\mid X)}+Y(0)\frac{\big{(}Y(0)-\mathbb{E}[Y(0)\mid X=x]\big{)}}{w(0\mid X)}\right\}+\frac{T}{S}\Big{\{}\theta^{2}(\widetilde{X};\beta_{0})\Big{\}}\right].$
Here, we used Lemma B.1.
@@
$T\mathrm{CR}\big{(}\theta\big{)}=\mathbb{E}_{W}\left[Y(1)\frac{\big{(}Y(1)-\mathbb{E}[Y(1)\mid X]\big{)}}{w(1\mid X)}+Y(0)\frac{\big{(}Y(0)-\mathbb{E}[Y(0)\mid X]\big{)}}{w(0\mid X)}\right]=\tau^{*}(\omega,p).$
B.5 Derivation of the Semiparametric Efficiency Bound
We extend the Cramér-Rao lower for a case with semiparametric models. As we did in Appendix B.4, we regard the whole $T+S$ data at hand as one observation and consider the case where we observe $m$ observations. Then, as $m$ goes to infinity, the total sample size $T^{\prime}:=Tm$ goes to infinity. Since we consider a case where $T+S$ samples are repeatedly observed from the same distribution independently, we can employ the conventional approach for deriving the semiparametric efficiency bound. Therefore, to employ this strategy, only in this section, we regard $T^{\prime}$ as the total “hypothetical” sample size in the asymptotic analysis regarding $m\to\infty$ while fixing $T$.
@@ B.5 Derivation of the Semiparametric Efficiency Bound
$\mathcal{T}\coloneqq\left\{\sum^{T}_{t=1}\left(\sum_{d\in\{1,0\}}\mathbbm{1}[A_{t}=d]\left\{s^{d}(Y_{t}\mid X_{t};\beta_{0})+\frac{\dot{w}(d\mid X_{t};\beta_{0})}{w(d\mid X_{t};\beta_{0})}\right\}+\zeta(X_{t};\beta_{0})\right)+\sum^{S}_{s=1}\widetilde{\zeta}\left(\widetilde{X}_{s};\beta_{0}\right)\in L_{2}(W)\right\}.$
where
@@
$\frac{\partial\theta(\beta_{0})}{\partial\beta}\coloneqq\mathbb{E}_{W}\Big{[}\psi(W)S(W;\beta_{0})\Big{]},$ (6)
@@
Proposition B.3 (From Theorem 25.20 in van der Vaart (1998).).
@@
This states that $\mathrm{Var}\big{(}\phi(W)\big{)}$ is a lower bound in estimating $\theta_{0}$. We call $T\mathrm{Var}\big{(}\phi(W)\big{)}$ the semiparametric efficiency bound because what we want to consider is a lower bound of $\sqrt{T}\left(\widehat{\theta}-\theta_{0}\right)$.
@@
$\psi(W)=\frac{1}{T}\sum^{T}_{t=1}\left\{\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X]\big{)}}{w(0\mid X_{t})}\right\}\frac{q(X_{t})}{p(X_{t})}+\frac{1}{S}\sum^{S}_{s=1}\Big{\{}\theta_{0}(\widetilde{X}_{s})-\theta_{0}\Big{\}}.$
The (scaled) efficiency bound $T\mathrm{Var}\big{(}\phi(W)\big{)}$ is
$\mathbb{E}_{X\sim p(x)}\left[\Bigg{\{}\frac{\sigma^{2}_{0}(1)(X)}{w(1\mid X)}+\frac{\sigma^{2}_{0}(0)(X)}{w(0\mid X)}\Bigg{\}}\frac{q^{2}(X)}{p^{2}(X)}\right]+\frac{T}{S}\mathrm{Var}_{\widetilde{X}\sim q(x)}\big{(}\theta(\widetilde{X})\big{)}.$
@@
As discussed in Appendices B.2, we define parametric submodels as
$p(W;\theta)$ $=\prod^{T}_{t=1}\prod_{d\in\{1,0\}}\Big{\{}r^{d}(Y_{t}\mid X_{t};\beta)w(d\mid X_{t};\beta)\Big{\}}^{A_{t}=d}p(X_{t};\beta)\prod^{S}_{s=1}q(\widetilde{X}_{s};\beta),$
@@
Then, recall that in Appendix B.2, we define the score function as
$S(W;\beta)\coloneqq\frac{\mathrm{d}}{\mathrm{d}\beta}\left\{\sum^{T}_{t=1}\log\big{(}f(Y_{t},A_{t},X_{t};\beta)+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\big{)}\right\}$
$=\sum^{T}_{t=1}\left\{\mathbbm{1}[A_{t}=1]\left\{s^{1}(Y_{t}\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right\}+\mathbbm{1}[A_{t}=0]\left\{s^{0}(Y_{t}\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right\}+\zeta(X_{t};\beta)\right\}$
@@
As shown in Lemma B.1, we have
@@
Then, based on the above preparation, we prove Lemma B.4 by guess and verify.
Proof of Lemma B.4.
Let us denote $\mathbb{E}_{X\sim q(x;),(Y,A)\sim f(y,a\mid X)}$ by $\mathbb{E}_{W}$. We first verify that $\mathbb{E}[\Psi^{2}(W)]<\infty$. Here, we have
$\mathbb{E}_{W}\left[\left\{\frac{\mathbbm{1}[A=1]\big{(}Y-\mathbb{E}[Y(1)\mid X]\big{)}}{w(1\mid X)}-\frac{\mathbbm{1}[A=0]\big{(}Y-\mathbb{E}[Y(0)\mid X]\big{)}}{w(0\mid X)}\right\}\frac{q(X)}{p(X)}\mid X\right]=0.$
Therefore,
$\mathbb{E}_{W}\left[\left(\left\{\frac{\mathbbm{1}[A=1]\big{(}Y-\mathbb{E}[Y(1)\mid X]\big{)}}{w(1\mid X)}-\frac{\mathbbm{1}[A=0]\big{(}Y-\mathbb{E}[Y(0)\mid X]\big{)}}{w(0\mid X)}\right\}\frac{q(X)}{p(X)}\right)\left(\theta_{0}(\widetilde{X})-\theta_{0}\right)\right]$
$=\mathbb{E}_{W}\left[\left(\theta_{0}(\widetilde{X})-\theta_{0}\right)\mathbb{E}_{W}\left[\left(\left\{\frac{\mathbbm{1}[A=1]\big{(}Y-\mathbb{E}[Y(1)\mid X]\big{)}}{w(1\mid X)}-\frac{\mathbbm{1}[A=0]\big{(}Y-\mathbb{E}[Y(0)\mid X]\big{)}}{w(0\mid X)}\right\}\frac{q(X)}{p(X)}\right)\mid X\right]\right]$
@@
$=\mathbb{E}_{W}\left[\left(\frac{1}{T}\sum^{T}_{t=1}\left\{\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X_{t}]\big{)}}{w(0\mid X_{t})}\right\}\frac{q(X_{t})}{p(X_{t})}+\frac{1}{S}\sum^{S}_{s=1}\Big{\{}\theta_{0}(\widetilde{X}_{s})-\theta_{0}\Big{\}}\right)^{2}\right]$
$=\mathbb{E}_{W}\left[\frac{1}{T}\Bigg{\{}\frac{\sigma^{2}_{0}(1)(X_{t})}{w(1\mid X_{t})}+\frac{\sigma^{2}_{0}(0)(X_{t})}{w(0\mid X_{t})}\Bigg{\}}\frac{q^{2}(X_{t})}{p^{2}(X_{t})}+\frac{1}{S}\Big{\{}\theta_{0}(\widetilde{X}_{s})-\theta_{0}\Big{\}}^{2}\right]<\infty.$
@@
$=\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\left\{\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X_{t}]\big{)}}{w(0\mid X_{t})}\right\}\frac{q(X_{t})}{p(X_{t})}+\frac{1}{S}\sum^{S}_{s=1}\Big{\{}\theta_{0}(\widetilde{X}_{s})-\theta_{0}\Big{\}}\Bigg{]}$
×$\times\left(\sum^{T}_{t=1}\left\{\mathbbm{1}[A_{t}=1]\left\{s^{1}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(1\mid X;\beta)}{w(1\mid X_{t};\beta)}\right\}+\mathbbm{1}[A_{t}=0]\left\{s^{0}(Y_{t}(0)\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right\}\right\}+\sum^{S}_{s=1}\widetilde{\zeta}(\widetilde{X}_{s};\beta)\right)\Bigg{]}$
$=\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A=1]\big{(}Y_{t}-\mathbb{E}[Y(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}\frac{q(X_{t})}{p(X_{t})}\left\{s^{1}(Y(1)\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right\}$

### v2

@@ A.2 Parametric Submodels
$f(y,a,x;\beta)=\prod_{d\in\{1,0\}}\left(r^{d}(y\mid x;\beta)w(d\mid x;\beta)\right)^{a=d}p(x;\beta)$ (4)
@@
We also denote $\prod_{d\in\{1,0\}}\left\{r^{d}(y\mid x;\beta)w(d\mid x;\beta)\right\}^{a=d}$ by $f(y,a\mid x;\beta)$.
@@
$\sum^{T}_{t=1}\log\left(f(Y_{t},A_{t},X_{t};\beta)\right)+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\big{)}.$
By taking the derivative of $\sum^{T}_{t=1}\log\left(f(Y_{t},A_{t},X_{t};\beta)\right)+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\big{)}$ with respect to $\beta$, we can obtain the corresponding score as
$S(W;\beta)\coloneqq\frac{\mathrm{d}}{\mathrm{d}\beta}\left(\sum^{T}_{t=1}\log\left(f(Y_{t},A_{t},X_{t};\beta)\right)+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\big{)}\right)$
$=\sum^{T}_{t=1}\left(\mathbbm{1}[A_{t}=1]\left\{s^{1}(Y_{t}\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right\}+\mathbbm{1}[A_{t}=0]\left(s^{0}(Y_{t}\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right)+\zeta(X_{t};\beta)\right)$
@@ A.3 ATE Function
$\theta(X;\beta)$ $\coloneqq\mathbb{E}_{Y(1)\sim r^{1}(y(1)\mid X;\beta)}[Y(1)\mid X]-\mathbb{E}_{Y(0)\sim r^{0}(y(0)\mid X;\beta)}[Y(0)\mid X].$
@@
A.4 Cramér-Rao Lower Bound
@@
$s^{1}(y\mid x;\beta)$ $=\frac{1}{T}\frac{\big{(}y-\mathbb{E}_{Y(1)\sim r^{d}(y(1)\mid x;\beta)}[Y(1)\mid X=x]\big{)}}{w(1\mid x)}$ (5)
$s^{0}(y\mid x;\beta)$ $=-\frac{1}{T}\frac{\big{(}y-\mathbb{E}_{Y(0)\sim r^{d}(y(0)\mid x;\beta)}[Y(0)\mid X=x]\big{)}}{w(0\mid x)}$
@@
$\zeta(x;\beta)$ $=0,$
$\widetilde{\zeta}(\widetilde{x};\beta)$ $=\frac{1}{S}\Big{\{}\theta(\widetilde{x};\beta)-\theta(\beta)\Big{\}}.$
@@
$\mathbb{E}_{W}\left[\left(\nabla_{\beta}\left(\sum^{T}_{t=1}\log\left(f(Y_{t},A_{t},X_{t};\beta)\right)+\sum^{S}_{s=1}\log\left(q\left(\widetilde{X}_{s};\beta\right)\right)\right)\right)^{2}\right]^{-1}.$
@@
$\mathrm{CR}(\theta)\coloneqq\nabla_{\beta}\theta(\beta)\mathbb{E}_{W}\left[\left(\nabla_{\beta}\left(\sum^{T}_{t=1}\log\left(f(Y_{t},A_{t},X_{t};\beta)\right)+\sum^{S}_{s=1}\log\left(q\left(\widetilde{X}_{s};\beta\right)\right)\right)\right)^{2}\right]^{-1}\nabla_{\beta}\theta(\beta).$
Theorem A.2.
@@
$C$ $\coloneqq\mathbb{E}_{W}\Big{[}Y(1)s^{1}(Y(1)\mid X;\beta_{0})-Y(0)s^{0}(Y(0)\mid X;\beta_{0})+\theta(\widetilde{X};\beta_{0})\widetilde{\zeta}(\widetilde{X};\beta_{0})\Big{]},$
$D$ $\coloneqq\mathbb{E}_{W}\left[\left(\nabla_{\beta}\left(\sum^{T}_{t=1}\log\big{(}f(Y_{t},A_{t},X_{t};\beta_{0})\big{)}+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta_{0})\right)\right)^{2}\right].$
Here, we used Lemma A.1.
@@
$T\mathrm{CR}\big{(}\theta\big{)}$
$=T\mathbb{E}_{W}\left[\frac{1}{T}\left(Y(1)\frac{\big{(}Y(1)-\mathbb{E}[Y(1)\mid X]\big{)}}{w(1\mid X)}+Y(0)\frac{\big{(}Y(0)-\mathbb{E}[Y(0)\mid X]\big{)}}{w(0\mid X)}\right)+\frac{1}{S}\theta(\widetilde{X};\beta_{0})\left(\theta(\widetilde{X};\beta_{0})-\theta(\beta_{0})\right)\right]$
×$\ \ \ \times\mathbb{E}_{W}\left[T\left(\nabla_{\beta}\log\big{(}f(Y,A,X;\beta_{0})\big{)}\right)^{2}+S\left(\nabla_{\beta}\log\big{(}q(\widetilde{X};\beta_{0})\right)^{2}\right]^{-1}$
$=\mathbb{E}_{W}\left[\left(\frac{\big{(}Y(1)-\mathbb{E}[Y(1)\mid X]\big{)}^{2}}{w(1\mid X)}+\frac{\big{(}Y(0)-\mathbb{E}[Y(0)\mid X]\big{)}^{2}}{w(0\mid X)}\right)+\frac{T}{S}\left(\theta^{2}(\widetilde{X};\beta_{0})\right)\right].$
@@
$T\mathrm{CR}\big{(}\theta\big{)}=\mathbb{E}_{W}\left[\left(\frac{\big{(}Y(1)-\mathbb{E}[Y(1)\mid X]\big{)}^{2}}{w(1\mid X)}+\frac{\big{(}Y(0)-\mathbb{E}[Y(0)\mid X]\big{)}^{2}}{w(0\mid X)}\right)+\frac{T}{S}\left(\theta^{2}(\widetilde{X};\beta_{0})\right)\right]=\tau^{*}(\omega,p).$
A.5 Derivation of the Semiparametric Efficiency Bound
We extend the Cramér-Rao lower for a case with semiparametric models. As we did in Appendix A.4, we regard the whole $T+S$ data at hand as one observation and consider the case where we observe $m$ observations. Then, as $m$ goes to infinity, the total sample size $T^{\prime}:=Tm$ goes to infinity. Since we consider a case where $T+S$ samples are repeatedly observed from the same distribution independently, we can employ the conventional approach for deriving the semiparametric efficiency bound. Therefore, to employ this strategy, only in this section, we regard $T^{\prime}$ as the total “hypothetical” sample size in the asymptotic analysis regarding $m\to\infty$ while fixing $T$.
@@ A.5 Derivation of the Semiparametric Efficiency Bound
$\mathcal{T}\coloneqq\left\{\sum^{T}_{t=1}\left(\sum_{d\in\{1,0\}}\mathbbm{1}[A_{t}=d]\left(s^{d}(Y_{t}\mid X_{t};\beta_{0})+\frac{\dot{w}(d\mid X_{t};\beta_{0})}{w(d\mid X_{t};\beta_{0})}\right)+\zeta(X_{t};\beta_{0})\right)+\sum^{S}_{s=1}\widetilde{\zeta}\left(\widetilde{X}_{s};\beta_{0}\right)\in L_{2}(W)\right\},$
where we have
@@
$\frac{\partial\theta(\beta_{0})}{\partial\beta}\coloneqq\mathbb{E}_{W}\Big{[}\psi(W)S(W;\beta_{0})\Big{]}.$ (6)
@@
Proposition A.3 (From Theorem 25.20 in van der Vaart (1998).).
@@
This states that $\mathrm{Var}\big{(}\psi(W)\big{)}$ is a lower bound in estimating $\theta_{0}$. We call $T\mathrm{Var}\big{(}\psi(W)\big{)}$ the semiparametric efficiency bound because what we want to consider is a lower bound of $\sqrt{T}\left(\widehat{\theta}-\theta_{0}\right)$.
@@
$\psi(W)=\frac{1}{T}\sum^{T}_{t=1}\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X]\big{)}}{w(0\mid X_{t})}\right)\frac{q(X_{t})}{p(X_{t})}+\frac{1}{S}\sum^{S}_{s=1}\left(\theta_{0}(\widetilde{X}_{s})-\theta_{0}\right).$
The (scaled) efficiency bound $T\mathrm{Var}\big{(}\psi(W)\big{)}$ is
$\mathbb{E}_{X\sim p(x)}\left[\left(\frac{\sigma^{2}_{0}(1)(X)}{w(1\mid X)}+\frac{\sigma^{2}_{0}(0)(X)}{w(0\mid X)}\right)\frac{q^{2}(X)}{p^{2}(X)}\right]+\frac{T}{S}\mathrm{Var}_{\widetilde{X}\sim q(x)}\big{(}\theta(\widetilde{X})\big{)}.$
@@
As discussed in Appendices A.2, we define parametric submodels as
$p(W;\beta)$ $=\prod^{T}_{t=1}\prod_{d\in\{1,0\}}\left(r^{d}(Y_{t}\mid X_{t};\beta)w(d\mid X_{t};\beta)\right)^{A_{t}=d}p(X_{t};\beta)\prod^{S}_{s=1}q(\widetilde{X}_{s};\beta),$
@@
Then, recall that in Appendix A.2, we define the score function as
$S(W;\beta)\coloneqq\frac{\mathrm{d}}{\mathrm{d}\beta}\left(\sum^{T}_{t=1}\log\left(f(Y_{t},A_{t},X_{t};\beta)\right)+\sum^{S}_{s=1}\log\big{(}q(\widetilde{X}_{s};\beta)\big{)}\right)$
$=\sum^{T}_{t=1}\left(\mathbbm{1}[A_{t}=1]\left(s^{1}(Y_{t}\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right)+\mathbbm{1}[A_{t}=0]\left(s^{0}(Y_{t}\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right)+\zeta(X_{t};\beta)\right)$
@@
As shown in Lemma A.1, we have
@@
Then, based on the above preparation, we prove Lemma A.4 by guess and verify.
Proof of Lemma A.4.
Let us denote $\mathbb{E}_{\widetilde{X}\sim q(x),(Y,A)\sim f(y,a\mid X)}$ by $\mathbb{E}_{W}$. We first verify that $\mathbb{E}_{W}[\psi^{2}(W)]<\infty$. Here, we have
$\mathbb{E}_{W}\left[\left(\frac{\mathbbm{1}[A=1]\big{(}Y-\mathbb{E}[Y(1)\mid\widetilde{X}]\big{)}}{w(1\mid\widetilde{X})}-\frac{\mathbbm{1}[A=0]\big{(}Y-\mathbb{E}[Y(0)\mid\widetilde{X}]\big{)}}{w(0\mid\widetilde{X})}\right)\frac{q(\widetilde{X})}{p(\widetilde{X})}\mid\widetilde{X}\right]=0.$
Therefore, we have
$\mathbb{E}_{W}\left[\left(\left(\frac{\mathbbm{1}[A=1]\big{(}Y-\mathbb{E}[Y(1)\mid\widetilde{X}]\big{)}}{w(1\mid\widetilde{X})}-\frac{\mathbbm{1}[A=0]\big{(}Y-\mathbb{E}[Y(0)\mid\widetilde{X}]\big{)}}{w(0\mid\widetilde{X})}\right)\frac{q(\widetilde{X})}{p(\widetilde{X})}\right)\left(\theta_{0}(\widetilde{X})-\theta_{0}\right)\right]$
$=\mathbb{E}_{W}\left[\left(\theta_{0}(\widetilde{X})-\theta_{0}\right)\mathbb{E}_{W}\left[\left(\frac{\mathbbm{1}[A=1]\big{(}Y-\mathbb{E}[Y(1)\mid\widetilde{X}]\big{)}}{w(1\mid\widetilde{X})}-\frac{\mathbbm{1}[A=0]\big{(}Y-\mathbb{E}[Y(0)\mid\widetilde{X}]\big{)}}{w(0\mid X)}\right)\frac{q(\widetilde{X})}{p(\widetilde{X})}\mid\widetilde{X}\right]\right]$
@@
$=\mathbb{E}_{W}\Bigg{[}\Bigg{(}\frac{1}{T}\sum^{T}_{t=1}\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid\widetilde{X}_{t}]\big{)}}{w(1\mid\widetilde{X}_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid\widetilde{X}_{t}]\big{)}}{w(0\mid\widetilde{X}_{t})}\right)\frac{q(\widetilde{X}_{t})}{p(\widetilde{X}_{t})}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{S}\sum^{S}_{s=1}\left(\theta_{0}(\widetilde{X}_{s})-\theta_{0}\right)\Bigg{)}^{2}\Bigg{]}$
$=\mathbb{E}_{W}\left[\frac{1}{T}\left(\frac{\sigma^{2}_{0}(1)(\widetilde{X}_{t})}{w(1\mid\widetilde{X}_{t})}+\frac{\sigma^{2}_{0}(0)\widetilde{X}_{t})}{w(0\mid\widetilde{X}_{t})}\right)\frac{q^{2}(X_{t})}{p^{2}(\widetilde{X}_{t})}+\frac{1}{S}\Big{\{}\theta_{0}(\widetilde{X}_{s})-\theta_{0}\Big{\}}^{2}\right]<\infty.$
@@
$=\mathbb{E}_{W}\Bigg{[}\Bigg{(}\frac{1}{T}\sum^{T}_{t=1}\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X_{t}]\big{)}}{w(0\mid X_{t})}\right)\frac{q(X_{t})}{p(X_{t})}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{S}\sum^{S}_{s=1}\Big{\{}\theta_{0}(\widetilde{X}_{s})-\theta_{0}\Big{\}}\Bigg{)}$
×$\times\Bigg{(}\sum^{T}_{t=1}\left(\mathbbm{1}[A_{t}=1]\left(s^{1}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(1\mid X;\beta)}{w(1\mid X_{t};\beta)}\right)+\mathbbm{1}[A_{t}=0]\left(s^{0}(Y_{t}(0)\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right)\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\sum^{S}_{s=1}\widetilde{\zeta}(\widetilde{X}_{s};\beta)\Bigg{)}\Bigg{]}$
$=\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}\frac{q(X_{t})}{p(X_{t})}\left(s^{1}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right)$